# Towards Tracing Trustworthiness Dynamics: Revisiting Pre-training Period of Large Language Models

### v1

@@ Appendix D Full Linear Probing Results
The full linear probing results from 360 checkpoints in five trustworthiness dimensions are shown in Figure 7,8,9,10,11. Overall, the experimental observations and conclusions are consistent with Section 2.3. Results from five datasets together suggest that middle layer representations exhibit linearly separable patterns. Furthermore, the probing accuracy increases during the initial phase of pre-training, followed by fluctuation throughout the remaining pre-training period.
@@ E.1 Mutual Information and HSIC
Mutual information is a measure of the mutual dependence between the two variables. However, because of the difficulty to accurately compute mutual information (Kraskov et al., 2004), we follow Ma et al. (2020) to use HSIC (Gretton et al., 2005) as an estimator of mutual information. HSIC (Gretton et al., 2005) also indicates the dependency between two random variables. For other kinds of estimation, please refer to Appendix E.3 in Zhou et al. (2023).
@@ E.2 Mutual Information Results across Five Trustworthiness Dimensions
Figure 12,13,14,15,16 show the trend of mutual information on five trustworthiness dimensions. The results are also consistent with the dynamics in Section 4.3. The phase transition from “fitting” to “compression” is also applicable: there are also two phases during pre-training. In the first and shorter phase, both $I(T,X)$ and $I(T,Y)$ increase. While in the second and much longer phase, $I(T,X)$ decreases and $I(T,Y)$ continues to increase. There are some fluctuations of $I(T,Y)$ for Toxigen, which may be due to the instability of pre-training.
@@
Appendix F Unlocking the Potential of Pre-trained Checkpoints through Proxy-tuning
The linear probe results of LLM360 and its evaluations across all checkpoints on TruthfulQA indicate that checkpoints during pre-training have already developed modeling capabilities for truthworthiness. Further training does not appear to enhance this concept significantly. However, cause the gap between latent space representation and model output Ravichander et al. (2021), strong representation seems not to be well applied. To address this, we attempt to shift the original predictions of the checkpoints during pre-training to enhance their utilization capabilities.
F.1 Proxy-Tuning to Checkpoints during Pre-training
Proxy-tuning applies the prediction differences between the tuned model and untuned model to shift the original predictions of a base model in the direction of tuning Liu et al. (2024); Mitchell et al. (2023). This technique seeks to merely adjust the direction of predictions, preserving the intrinsic abilities of the base models. Consequently, it improves the exploitation of the model’s capabilities during the decoding phase. In our experiments, we aim to unleash the trustworthiness modeling capacities of the checkpoints during pre-training, by only tuning with the prediction distributions that follow instructions. Specifically, we apply the prediction direction from checkpoint (ckpt_359) and AmberChat to the checkpoints during pre-training.
F.2 Performance Enhancement on TruthfulQA via Proxy-Tuning
@@
Appendix G Cases of TruthfulQA Answers under Different Perplexity

### v2

@@ Appendix D Full Linear Probing Results
The full linear probing results from 360 checkpoints in five trustworthiness dimensions are shown in Figure 7,8,9,10,11. Overall, the experimental observations and conclusions are consistent with Section 2.3. Results from five datasets together suggest that middle-layer representations exhibit linearly separable patterns. Furthermore, the probing accuracy increases during the initial phase of pre-training, followed by fluctuation throughout the remaining pre-training period.
@@ E.1 Mutual Information and HSIC
Mutual information is a measure of the mutual dependence between the two variables. However, because of the difficulty to accurately compute mutual information (Kraskov et al., 2004), we follow Ma et al. (2020) to use HSIC (Gretton et al., 2005) as an estimator of mutual information. HSIC (Gretton et al., 2005) also indicates the dependency between two random variables. For other kinds of estimation, please refer to Appendix E.3 in Zhou et al. (2023b).
@@ E.2 Mutual Information Results across Five Trustworthiness Dimensions
Figure 12,13,14,15,16 show the trend of mutual information on five trustworthiness dimensions. The results are also consistent with the dynamics in Section 4.3. The phase transition from “fitting” to “compression” is also applicable: there are also two phases during pre-training. In the first and shorter phase, both $I(T,X)$ and $I(T,Y)$ increase. While in the second and much longer phase, $I(T,X)$ decreases, and $I(T,Y)$ continues to increase. There are some fluctuations of $I(T,Y)$ for Toxigen, which may be due to the instability of pre-training.
@@
Appendix F Experimental Results on Another Series of LLMs’ Checkpoints
To further demonstrate the generalization performance of the observations in this work, we conduct additional experiments on a recently released open-source model named OLMo Groeneveld et al. (2024). OLMo provides all intermediate checkpoints during the pre-training period. Additionally, OLMo also releases an instruction-tuned model named OLMo-7B-SFT and an aligned model through DPO named OLMo-7B-Instruct. Currently, the largest model size available in the OLMo project is 7B parameters.
Figure 17: The linear probe accuracy on five trustworthiness dimensions for the first 60 pre-training checkpoints of OLMo. For each checkpoint, we report the results from layers {0, 6, 12, 18, 24, 30}.
We follow the experimental settings in Section 2.2 to conduct probing experiments on OLMo. The results are shown in Figure 17. Figure 17 demonstrates that after the early pre-training period, middle layer representations of LLMs have already developed linearly separable patterns about trustworthiness. This aligns with the results obtained from LLM360, as introduced in Section 2.3.
Table 4: Results of activation intervention on OLMo in TruthfulQA and StereoSet. Format and significance markers remain consistent with Table 1. $\bm{v}_{ckpt\_{279}}$ denotes the steering vector extracted from the $279$-th checkpoint.
Model TruthfulQA Metrics StereoSet Metric
[…]
OLMo-7B-SFT 0.4668 0.9803 0.4576 0.5471
OLMo-7B-SFT + $\bm{v}_{ckpt\_{279}}$ 0.6708 0.9631 0.6460 0.5789
We further conduct activation intervention experiments on the TruthfulQA and StereoSet datasets, following the experimental settings in Section 3.2. The results of the activation intervention on TruthfulQA and StereoSet datasets are presented separately in Table 4. We observe that steering vectors $V_{ckpt\_279}$ derived from pre-training checkpoints could improve the SFT model’s performance. This verifies steering vectors extracted from pre-training checkpoints could promisingly enhance the SFT model’s trustworthiness, and the experimental observation is consistent with the Observation 1 in Section 3.3.
Appendix G Unlocking the Potential of Pre-trained Checkpoints through Proxy-tuning
The linear probe results of LLM360 and its evaluations across all checkpoints on TruthfulQA indicate that checkpoints during pre-training have already developed modeling capabilities for truthworthiness. Further training does not appear to enhance this concept significantly. However, cause of the gap between latent space representation and model output Ravichander et al. (2021); strong representation seems not to be well applied. To address this, we attempt to shift the original predictions of the checkpoints during pre-training to enhance their utilization capabilities.
G.1 Proxy-Tuning to Checkpoints during Pre-training
Proxy-tuning applies the prediction differences between the tuned model and the untuned model to shift the original predictions of a base model in the direction of tuning Liu et al. (2024); Mitchell et al. (2023). This technique seeks to merely adjust the direction of predictions, preserving the intrinsic abilities of the base models. Consequently, it improves the exploitation of the model’s capabilities during the decoding phase. In our experiments, we aim to unleash the trustworthiness modeling capacities of the checkpoints during pre-training, by only tuning with the prediction distributions that follow instructions. Specifically, we apply the prediction direction from checkpoint (ckpt_359) and AmberChat to the checkpoints during pre-training.
G.2 Performance Enhancement on TruthfulQA via Proxy-Tuning
@@
Appendix H Cases of TruthfulQA Answers under Different Perplexity